OpenStudy (anonymous):

In a survey of 1000 randomly chosen adults, 605 said that they used email. Calculate a 90% confidence interval for the proportion of adults in the whole population who use email.

OpenStudy (amistre64):

605/1000 is our "mean" as i recall

OpenStudy (amistre64):

have to review a little to make sure what to do next :)

OpenStudy (amistre64):

z*sqrt(pq/n)

OpenStudy (amistre64):

the zscore for 90% is ...

OpenStudy (amistre64):

\[\frac{605}{1000}\pm z*\sqrt{\frac{605}{1000}\frac{395}{1000}\frac{1}{1000}}\]

OpenStudy (anonymous):

Is it OK if you can find out some more background info on this matter? Like, maybe a site to explain this more?

OpenStudy (amistre64):

http://www.stat.wmich.edu/s160/book/node47.html this is what refreshed it for me

OpenStudy (anonymous):

Thanks for it... So, from the last step you wrote, how do you continue?

OpenStudy (amistre64):

you determine the zscore for 90% and plug it into the rest to determine your interval

OpenStudy (anonymous):

So, how do you determine the zscore? With a table... or?

OpenStudy (amistre64):

table, yes; or ti83

OpenStudy (anonymous):

I don't have a graphic calc... Would you use the Modulus Function table... or?

OpenStudy (amistre64):

oy this thing is slooowwww this morning

OpenStudy (amistre64):

this measures the z from the mean itself; so we want .4500 from the mean so that the total area from the mean is 90%

OpenStudy (amistre64):

.4500 in the feild is there in the middle; it rests neatly between 1.64 and 1.65 1.645 should be the zscore then

OpenStudy (amistre64):

\[.605\pm 1.645\sqrt{.605*.305*.001}\]

OpenStudy (amistre64):

the ;ower and upper limits of that stuff define the interval

OpenStudy (anonymous):

Hmm.. Ok... I don't have this table in my statistics book. Do you think they'll give it me in the exam?

OpenStudy (amistre64):

lol, i doubt it. the table should be on the back cover

OpenStudy (amistre64):

or in a leaflet; its called a ztable

OpenStudy (anonymous):

Hmmm... I hope they do! ztable? What's the difference between a ztable and a modulus function table?

OpenStudy (amistre64):

depends on what a modulus function table is :)

OpenStudy (anonymous):

I mean a normal distribution table...?

OpenStudy (anonymous):

But... the z table you first gave looks different than the normal distribution table?

OpenStudy (amistre64):

there are a few ways that it is written up depending on how the authors like to display it the main difference is in what they measure from.

OpenStudy (amistre64):

in my book; they measure from the mean itself; in other books they measure from a tail

OpenStudy (amistre64):